OpenStudy (perii224):

PLEASE HELP WILL MEDAL Given f(x) = x + 9 and g(x) = 5x, what is f(3) + g(-3)? -3 3 -27 27 Using f(x) = 3x - 3 and g(x) = 2x, what is f(4) - g(2) ? 5 -5 10 -10 Given f(x) = 6x + 2 and g(x) = x3 , what is f(x) g(x)? 6x4 + 2 8x3 6x4 + 2x3 6x3 + 2

OpenStudy (michele_laino):

hint: question #1 we have \(f(3)=3+9=...\), and \(g(-3)=5 \cdot (-3)=...\)